OpenStudy (anonymous):

y=-3x+1 y=-x+5 solve by graphing

OpenStudy (anonymous):

\[3x+y=1\] ------------i) \[x+y=5\] ------------ii) \[3x+y=1\] --------i)*1 \[3x+3y=15\] --------ii)*3 (-) (-) (-) ----------------- 0 -2y= -14 then y=7 and x+7=5 x=-7+5 x=-2 x=-2 , y=7

OpenStudy (anonymous):

sorry i can't can prove this graph in this screen i proved that in my copy book if you need urgent i will sent the attachment copy